If f(x)=x^2-1 and g(x)=x+1, find f(g(x)) and g(f(x))
Answer by Cromlix(4381)

Hi there,
f(x) = x^2 - 1
g(x) = x + 1
.............
f(g(x)) = (x + 1)^2 - 1
f(g(x)) = x^2 + 2x
..........
g(f(x)) = (x^2 - 1) + 1
g(f(x)) = x^2
..............
Hope this helps :-)